# Nonparametric bounds for vaccine effects in randomized trials

### v1

@@ I.1 Proof that the LP-based bounds in Proposition 5 are sharp
with no dependency on $S$. Following this result, the one-sided monotonicity-based bounds under Figures 3(a)3(d) coincide with those under Figure 2. Consequently, adding Assumption 8 yields the same two-sided monotonicity-based bounds as in Figure 2.
Let $\gamma^{IV}_{s.a}=Pr_{IV}(S=s\mid A=a)$. Under Figures 3(a)3(d) and the assumption that $Y\perp M\mid A,B,S,U$, it follows that $E(Y^{a,m=1})$ for $a\in\{0,1\}$ is identified by
$\begin{split}E&(Y^{a,m=1})\\ &=\sum_{s}E_{IV}\!\left(Y\mid A=a,M=1,B=1,S=s\right)\gamma^{IV}_{s.a}\\ &=\sum_{s,u}E_{IV}\!\left(Y\mid A=a,M=1,B=1,S=s,U=u\right)\Pr_{IV}(U=u\mid A=a,M=1,B=1,S=s)\gamma^{IV}_{s.a}\\ &=\sum_{s,u}E_{IV}\!\left(Y\mid A=a,M=1,B=1,S=s,U=u\right)\Pr_{IV}(U=u)\gamma^{IV}_{s.a}\\ &=\sum_{s,u}E\!\left(Y\mid A=a,M=-1,B=1,S=s,U=u\right)\Pr(U=u)\gamma^{IV}_{s.a}\\ &=E_{U}\!\left[\sum_{s}E\!\left(Y\mid A=a,B=1,S=s,U\right)\gamma_{s.a}\right]\\ &=\sum_{s}E_{U}\!\left[E\!\left(Y\mid A=a,B=1,S=s,U\right)\gamma_{s.a}\right].\end{split}$ (40)
@@
$\begin{split}E&(Y\mid A=a,B=1)\\ &=\sum_{s,u}E\!\left(Y\mid A=a,B=1,S=s,U=u\right)\gamma_{s.a}\Pr(U=u\mid A=a,B=1,S=s)\\ &=\sum_{s,u}E\!\left(Y\mid A=a,B=1,S=s,U=u\right)\gamma_{s.a}\frac{\Pr(B=1\mid A=a,U=u,S=s)\Pr(U=u\mid A=a,S=s)}{\Pr(B=1\mid A=a,S=s)}\\ &=\sum_{s,u}E\!\left(Y\mid A=a,B=1,S=s,U=u\right)\gamma_{s.a}\Pr(U=u)\frac{\Pr(B=1\mid A=a,U=u,S=s)}{\Pr(B=1\mid A=a,S=s)}\\ &=E_{U}\!\left[\sum_{s}E\!\left(Y\mid A=a,B=1,S=s,U\right)\gamma_{s.a}\frac{\Pr(B=1\mid A=a,S=s,U)}{\Pr(B=1\mid A=a,S=s)}\right]\\ &=\sum_{s}E_{U}\!\left[E\!\left(Y\mid A=a,B=1,S=s,U\right)\gamma_{s.a}\frac{\Pr(B=1\mid A=a,S=s,U)}{\Pr(B=1\mid A=a,S=s)}\right].\end{split}$ (41)
Let $\nu(s,u)=\frac{\Pr(B=1\mid A=a,S=s,U=u)}{\Pr(B=1\mid A=a,S=s)}$. Then $\nu(s,u)$ is nondecreasing in $u$ since $\Pr(B=1\mid A=a,S=s,U=u)$ is nondecreasing in $u$, and $E_{U}[\nu(s,U)]=1$. Thus, by Lemma 1 we have
$\begin{split}&\sum_{s}E_{U}\!\left[E\!\left(Y\mid A=a,B=1,S=s,U\right)\gamma_{s.a}\nu(s,U)\right]\\ &\;-\sum_{s}E_{U}\!\left[E\!\left(Y\mid A=a,B=1,S=s,U\right)\gamma_{s.a}\right]\\ &=\sum_{s}\operatorname{Cov}_{U}\!\left(E\!\left(Y\mid A=a,B=1,S=s,U\right)\gamma_{s.a},\,\nu(s,U)\right)\geq 0.\end{split}$ (42)
By substituting Equations 4041 into Equation LABEL:eq:cov2, we obtain
$E(Y\mid A=a,B=1)$ $=\sum_{s}E_{U}\!\left[E\!\left(Y\mid A=a,B=1,S=s,U\right)\gamma_{s.a}\nu(s,U)\right]$
$\geq\sum_{s}E_{U}\!\left[E\!\left(Y\mid A=a,B=1,S=s,U\right)\gamma_{s.a}\right]=E(Y^{a,m=1}).$
By Lemma 2, we can similarly prove the bounds for $E(Y^{a,m=0})$ and for the discordant-sign case.
L.1 Proof that the monotonicity-based bounds in Proposition 7 are sharp
Since:
@@
    Figure 2 is a special case of Figures 3(b)3(d), obtained by setting to zero some of the arrows $S\rightarrow Y$, $B\rightarrow Y$, and $U\rightarrow S$ in Figures 3(b)3(d); and
@@
    the analytical bounds under these figures coincide with those derived for Figure 2,
it follows that, because the bounds under Figure 2 are sharp, the bounds under Figures 3(b)3(d) are sharp as well.
@@
In this Appendix, we show that although Figure M.1 includes two common causes, it is still equivalent to Figures 3(b)3(d) in the sense that it suffices to condition on only one of the $U$s, such that $Y\perp_{IV}M\mid A,S,B,U_{i}$ for $i\in\{1,2\}$. Therefore, the bounds in Propositions 6 and 7 remain valid under Figure M.1.

### v2

@@ Appendix I Proof of Proposition 3: LP-based bounds under Figure 3(a)
with no dependency on $S$. Following this result, the one-sided monotonicity-based bounds under Figures 3(b)3(d) coincide with those under Figure 2. Consequently, adding Assumption 8 yields the same two-sided monotonicity-based bounds as in Figure 2. We will prove the result under Figure 3(d). Since Figures 3(b)3(c) are special cases of Figure 3(d), the result holds for them as well. In Section L.1, we recall a standard characterization of stochastic dominance; in Section L.2, we apply this result to show that $\gamma_{s.a1}\geq_{st}\gamma_{s.a}$ under a monotonicity condition; and, in Section L.3, we provide the proof that $E(Y\mid A=a,B=1)\geq E(Y^{a,m=1})$. Finally, in Section L.4, we present the assumption yielding reversed bounds.
In this section we use the notation $X\geq_{st}Y$ to denote that $X$ stochastically larger than $Y$.
L.1 Shaked and Shanthikumar, 2007, Lemma (1.A.7), pg. 4
Lemma 3.
Let $X$ and $Y$ be random variables. Then
$X\geq_{st}Y$
if and only if for every increasing function $\phi$ for which the expectations exist,
$E[\phi(X)]\geq E[\phi(Y)].$
L.2 Proof that $\gamma_{s.a1}\geq_{st}\gamma_{s.a}$
Lemma 4.
Assume that for any $a\in\{0,1\}$
$\Pr(B=1\mid A=a,S=1)\;\geq\;\Pr(B=1\mid A=a,S=0).$ (34)
Then, the distribution of $S$ given $(A=a,B=1)$ stochastically dominates the distribution of $S$ given $A=a$, i.e.
$\gamma_{s.a1}\geq_{st}\gamma_{s.a}.$
Proof.
Since $S\in\{0,1\}$, first-order stochastic dominance reduces to verifying
$\Pr(S=1\mid A=a,B=1)\;\geq\;\Pr(S=1\mid A=a).$ (35)
Let $p=\Pr(S=1\mid A=a)$ and write $w_{s}=\Pr(B=1\mid A=a,S=s)$ for $s=0,1$. By Bayes’ rule,
$\Pr(S=1\mid A=a,B=1)=\frac{w_{1}p}{\Pr(B=1\mid A=a)}.$
The law of total probability gives
$\Pr(B=1\mid A=a)=w_{1}p+w_{0}(1-p).$
Substituting yields
$\Pr(S=1\mid A=a,B=1)=\frac{w_{1}p}{w_{1}p+w_{0}(1-p)}.$
To show (35), note that
$\frac{w_{1}p}{w_{1}p+w_{0}(1-p)}\geq p\quad\Longleftrightarrow\quad p(1-p)(w_{1}-w_{0})\geq 0,$
which holds because $0\leq p\leq 1$ and $w_{1}\geq w_{0}$ by assumption (34). Thus $\gamma_{s.a1}\geq_{st}\gamma_{s.a}$.
L.3 Proof that $E(Y\mid A=a,B=1)\geq E(Y^{a,m=1})$
Let $\gamma^{IV}_{s.a}=Pr_{IV}(S=s\mid A=a)$. Under Figure 3(d) and the assumptions (1) $Y\perp M\mid_{VI}A,B,S,U$, (2) $\gamma^{IV}_{s.a}=\gamma_{s.a}$, and (3) $\Pr_{IV}(U=u|S=s,A=a)=\Pr(U=u|S=s,A=a)$, it follows that $E(Y^{a,m=1})$ for $a\in\{0,1\}$ is identified by
$\begin{split}E&(Y^{a,m=1})\\ &=\sum_{s}E_{IV}\!\left(Y\mid A=a,M=1,B=1,S=s\right)\gamma^{IV}_{s.a}\\ &=\sum_{s,u}E_{IV}\!\left(Y\mid A=a,M=1,B=1,S=s,U=u\right)\Pr_{IV}(U=u\mid A=a,M=1,B=1,S=s)\gamma^{IV}_{s.a}\\ &=\sum_{s,u}E_{IV}\!\left(Y\mid A=a,M=1,B=1,S=s,U=u\right)\Pr_{IV}(U=u|{S=s,A=a})\gamma^{IV}_{s.a}\\ &=\sum_{s,u}E\!\left(Y\mid A=a,M=-1,B=1,S=s,U=u\right)\Pr(U=u|{S=s,A=a})\gamma^{IV}_{s.a}\\ &=\sum_{s}E_{U}\!\left[E\!\left(Y\mid A=a,B=1,S=s,U\right)\gamma_{s.a}|{S=s,A=a}\right].\end{split}$ (36)
@@
$\begin{split}E&(Y\mid A=a,B=1)\\ &=\sum_{s}E\!\left(Y\mid A=a,B=1,S=s\right)\gamma_{s.a{b}}\\ &=\sum_{s,u}E\!\left(Y\mid A=a,B=1,S=s,U=u\right)\gamma_{s.a{b}}\Pr(U=u\mid A=a,B=1,S=s)\\ &=\sum_{s,u}E\!\left(Y\mid A=a,B=1,S=s,U=u\right)\gamma_{s.a{b}}\frac{\Pr(B=1\mid A=a,U=u,S=s)\Pr(U=u\mid A=a,S=s)}{\Pr(B=1\mid A=a,S=s)}\\ &=\sum_{s,u}E\!\left(Y\mid A=a,B=1,S=s,U=u\right)\gamma_{s.a{b}}\Pr(U=u|{A=a,S=s})\frac{\Pr(B=1\mid A=a,U=u,S=s)}{\Pr(B=1\mid A=a,S=s)}\\ &=\sum_{s}E_{U}\!\left[E\!\left(Y\mid A=a,B=1,S=s,U\right)\gamma_{s.a{b}}\frac{\Pr(B=1\mid A=a,S=s,U)}{\Pr(B=1\mid A=a,S=s)}\bigg|{A=a,S=s}\right].\end{split}$ (37)
Now, we are going to prove that $E(Y\mid A=a,B=1)-E(Y^{a,m=1})\geq 0$. Let $\nu_{s}(u)=\frac{\Pr(B=1\mid A=a,S=s,U=u)}{\Pr(B=1\mid A=a,S=s)}$ and $g_{s}(U)=E\!\left(Y\mid A=a,B=1,S=s,U\right)$. Assume that (1) $\Pr(B=1\mid A=a,S=s,U=u)$ and $E\!\left(Y\mid A=a,B=1,S=s,U\right)$ are both nondecreasing in $u$ for all $a$ and $s$, (2) $Pr(B=1|A=a,S=1)\geq Pr(B=1|A=a,S=0)$ and (3) $E_{U}(g_{s}(U)|A=a,S=s)=E(Y^{a,m}|A=a,S=s)$ is nondecreasing in $s$. Then, $E(Y\mid A=a,B=1)-E(Y^{a,m=1})$ can be written as
$\begin{split}E&(Y\mid A=a,B=1)-E(Y^{a,m=1})\\ &\overset{\mathrm{(1)}}{=}\sum_{s}E_{U}\!\left[{\gamma_{s.a1}}g_{s}(U)\nu_{s}(U)|A=a,S=s\right]-\sum_{s}E_{U}\!\left[{\gamma_{s.a}}g_{s}(U)|A=a,S=s\right]\\ &\overset{\mathrm{}}{=}\sum_{s}\bigg[[E_{U}\!\left[{\gamma_{s.a1}}g_{s}(U)\nu_{s}(U)|A=a,S=s\right]-E_{U}\!\left[{\gamma_{s.a}}g_{s}(U)|A=a,S=s\right]\bigg]\\ &\overset{\mathrm{(2)}}{=}\sum_{s}\bigg[{\gamma_{s.a1}}\operatorname{Cov}_{U}\!\left(g_{s}(U),\,\nu_{s}(U)|A=a,S=s\right)+(\gamma_{s.a1}-\gamma_{s.a})E_{U}(g_{s}(U)|A=a,S=s)\bigg]\\ &\overset{\mathrm{}}{=}E_{s.a1}\bigg[\operatorname{Cov}_{U}\!\left(g_{s}(U),\,\nu_{s}(U)|A=a,S=s\right)\bigg]\\ &+\big(E_{\gamma_{s.a1}}E_{U}(g_{s}(U)|A=a,S=s)-E_{\gamma_{s.a}}E_{U}(g_{s}(U)|A=a,S=s)\big)\\ &\overset{\mathrm{(3)}}{\geq}\ 0\end{split}$ (38)
Step (1) follows from the derivations above and the definitions of $\nu_{s}(u)$ and $g_{s}(U)$. Step (2) follows since $E_{U}[\nu_{s}(U)|A=a,S=s]=1$ and since it is easy to show that for any random variables $X,Y$, and constants $a,b$, E(aXY|C)E(bX|C)E(Y|C)=aCov(X,Y|C)+(ab)E(X|C)E(Y|C)E(aXY|C)-E(bX|C)E(Y|C)=aCov(X,Y|C)+(a-b)E(X|C)E(Y|C) where $C=\{A=a,S=s\}$. Step (3) follows from the following two points
  1. 1.
    We assume that $g_{s}(U)$ is nondecreasing in $u$ and that $\Pr(B=1\mid A=a,S=s,U=u)$ is nondecreasing in $u$ and thus $\nu_{s}(u)$ is nondecreasing in $u$. From Lemma 1 it follows that $\operatorname{Cov}_{U}\!\left(g_{s}(U),\,\nu_{s}(U)\right|A=a,S=s)\geq 0$.
  2. 2.
    We assume that $E_{U}(g_{s}(U)|A=a,S=s)$ is nondecreasing in $s$. By Lemma 4 we have that $\gamma_{s.a1}\geq_{st}\gamma_{s.a}$. Thus, by Lemma 3 we have that $\big(E_{\gamma_{s.a1}}E_{U}(g_{s}(U)|A=a,S=s)-E_{\gamma_{s.a}}E_{U}(g_{s}(U)|A=a,S=s)\big)\geq 0$.
L.4 Assumption guaranteeing reversed bounds
Assumption 14.
($U|S$ monotonicity) Consider the setting of Figures 3(b)3(d). Suppose that $Y\perp_{\mathcal{T}_{VI}}M|A,B,S,U$ the following conditions hold:
@@
    For all $a$ and $s$, exactly one of $\Pr\!\left(Y=1\mid A=a,B=1,S=s,U=u\right)$ and $\Pr(B=1\mid A=a,S=s,U=u)$ is nondecreasing in $u$, while the other is nonincreasing in $u$.
  3. (iii)
    For all $a$, $\Pr(B=1\mid A=a,S=s)$ is nondecreasing in $s$.
  4. (iv)
    For all $a$, $E(Y^{a,m}\mid A=a,S=s)$ is nonincreasing in $s$.
@@
In this Appendix, we show that although Figure M.1 includes two common causes, it is still equivalent to Figures 3(b)3(d) in the sense that it suffices to condition on only one of the $U$s, such that $Y\perp_{IV}M\mid A,S,B,U_{i}$ for $i\in\{1,2\}$. Therefore, the bounds in Propositions 5 and 6 remain valid under Figure M.1.